Q1. Differentiate y = 12x5 + 3x4 + 7x3 + x2 

Q2. Find the distance between the points (2, 3) and (0, 6)

Q3. The compound interest and simple interest on a certain sum for 2 years is N1230 and N1200 respectively. The rate of interest is same for both compound interest and simple interest and it is compounded annually. What is the principal?


1
Expert's answer
2021-05-02T09:40:40-0400

Solution. Q1 Using the rules of differentiation and the table of derivatives, we get


"y'=12\times5\times x^4+3\times4\times x^3+7\times3\times x^2+2\times x"

"y'=60x^4+12x^3+21x^2+2x"

Q2 We use the formula for the distance between two points with coordinates (x1,y1), (x2,y2)


"d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}."

Therefore, the distance between the points (2, 3) and (0, 6) is equal to


"d=\sqrt{(0-2)^2+(6-3)^2}=\sqrt{4+9}=\sqrt{13}"

Q3 Using Simple Interest Equation (Principal + Interest)


"A=P(1+it)"

where A is the total accrued amount (principal + interest), P is the principal amount, i is the rate of interest per year in decimal, t is time period involved in years.

The compound interest formula is as follows:


"T=P(1+i)^t"

where T is total accrued, including interest; P is the principal amount; i is the rate of interest per year in decimal, t is time period involved in years.

According to the condition of the problem t= 2years, T=1230, A=1200, and the rate of interest is the same for both compound interest and simple interest and it is compounded annually.

As result get


"P(1+2i+i^2)=1230"

"P(1+2i)=1200"

Subtracting the second from the first equation, we obtain


"Pi^2=30."

Hence


"P=\frac{30}{i^2}"

Substituting the resulting expression into the second equation, we get


"\frac{30}{i^2}(1+2i)=1200."

Therefore


"1+2i=40i^2"

"40i^2-2i-1=0"

"D=(-2)^2-4\times40\times(-1)=4+160=164"

"i_1=\frac{2-\sqrt{164}}{80}<0"

"i_2=\frac{2+\sqrt{164}}{80}"

As result the principal is equl to


"P=\frac{30}{i^2}=\frac{30}{(\frac{2+\sqrt{164}}{80})^2}\approx 876"


Answer. Q1


"y'=60x^4+12x^3+21x^2+2x"

Q2


"d=\sqrt{13}"

Q3


"P= 876"